OpenStudy (anonymous):

Find an equation of the tangent line to the hyperbola: x^(2)-2 y^(2) = 1 at the point (a,b).

OpenStudy (anonymous):

Here: http://openstudy.com/study#/updates/4f7ef8c8e4b0bfe8930b75a2

OpenStudy (anonymous):

i got slope = x/2y

OpenStudy (anonymous):

It should help you figure out the answer!

OpenStudy (anonymous):

is it y - 2b = (a/2b) (x - a)

OpenStudy (anonymous):

is this right or wrong

myininaya (myininaya):

Isn't the point on the line suppose to be (a,b) not (a,2b)?

myininaya (myininaya):

Make that correction and you are good.

OpenStudy (anonymous):

it should be y - b ?

myininaya (myininaya):

yep since the point is (a,b) not (a,2b) y-y1=m(x-x1) you found the slopt to be m=a/(2b) but this doesn't change the point from (a,b) to (a,2b) It is still (a,b) y-b=a/(2b)*(x-a)